OpenStudy (anonymous):

Let f(x)=7-2x^2. Find the value of f(4+h).

OpenStudy (anonymous):

\[ f(x)=7-2x^2.\] \[f(x)=7-2(4+h)^2.\]

OpenStudy (anonymous):

I know that but it has to be solved.

OpenStudy (anonymous):

by "solved" i take it you mean multiply out and combine like terms right?

OpenStudy (anonymous):

Yes, trying to work it out right now.

OpenStudy (anonymous):

I typed it exactly as it says on the paper.

OpenStudy (anonymous):

\[f(x)=7-2(4+h)^2\\ =7-2(4+h)(4+h)\] is a start

OpenStudy (anonymous):

leave the -2 out front \[7-2(16+8h+h^2)\] is the next step

OpenStudy (anonymous):

then distribute the \(-2\)

OpenStudy (anonymous):

So you foil after squaring? That is the part I always get confused at.

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

squaring means to multiply by itself so \[(a+b)^2=(a+b)(a+b)\] and \[(4+h)^2=(4+h)(4+h)\]

OpenStudy (anonymous):

Yeah, I think I got it

OpenStudy (anonymous):

it is always the case that \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

Answer is -25-16h-2h^2

OpenStudy (anonymous):

ok if you say so

OpenStudy (anonymous):

Or better written as -2h^2-16h-15

OpenStudy (anonymous):

yeah looks good to me

OpenStudy (anonymous):

btw i see in your profile that you are majoring in education do me a huge favor please do not every tell students to "solve" unless you are solving an equation what you are doing here is not solving, it is writing in standard form and don't even get me started on "foil"

OpenStudy (anonymous):

LOL. majoring in early childhood/special education. Likely hood of needing to do/teach this is slim.